OpenStudy (anonymous):

A particle is moving with the given data. Find the position of the particle. a(t) = cos t + sin t, s(0) = 1, v(0) = 4

OpenStudy (anonymous):

v(0)=-1 s(0)=1

OpenStudy (anonymous):

sooo s(t) would be 1?

OpenStudy (anonymous):

s(t)=-cost-sint

OpenStudy (anonymous):

thank you thank you thank you thank youuuu!!!

OpenStudy (anonymous):

sorry s(0)=-1

OpenStudy (anonymous):

You have to integrate the function a(t) twice to get it's position function. First we get the velocity function: v(t)=∫(cost+sint)dt=sint−cost+C Where C is a constant. Since v(0)=4 we get: v(0)=sin0−cos0+C=4 -1+C=4 and C=5 Now we integrate the velocity function to get the position function s(t): s(t)=∫(sint−cost+5)dt=−cost−sint+5t+D Where D is a constant. We have s(0)=1. So, s(0)=5(0)−sin(0)−cos(0)+D=1 0-0-1+D=1 D=2 So you position function is: s(t)=5t-sint-cost+2

OpenStudy (anonymous):

Thank you!